OpenStudy (anonymous):

Determine algebraically whether the function is even, odd, or neither even nor odd. f(x) = -9 x3 + 8x @ganeshie8

OpenStudy (anonymous):

odd, correct?

ganeshie8 (ganeshie8):

f(-x) = -9(-x)^2=3 + 8(-x) = 9x^3 - 8x = -(-9x^3 + 8x) = -f(x) => odd

ganeshie8 (ganeshie8):

good job !